OpenStudy (anonymous):

shell method bounded by x^2, y=8-7x, x=0

OpenStudy (aum):

rotated about y-axis?

OpenStudy (anonymous):

Yes

OpenStudy (aum):

Find where y = x^2 and y = 8 - 7x intersect to find the limits of integration.

OpenStudy (anonymous):

it looks like 1

OpenStudy (aum):

\[V = \int\limits_{0}^{1}2\pi x * (8-7x - x^2)dx\]

OpenStudy (anonymous):

so just substitute the upper and lower limits?

OpenStudy (anonymous):

after moving 2pi out?

OpenStudy (aum):

Multiply (8-7x-x^2) by x. Then integrate, term by term. Then substitute the upper and lower limits.

OpenStudy (anonymous):

I got: 4-(7/3)-(1/4) after multiplying and substituting

OpenStudy (aum):

Simplify the fractions. Don't forget the 2*pi outside.

OpenStudy (anonymous):

2pi*2/3

OpenStudy (anonymous):

4/3pi

OpenStudy (aum):

\( \Large 2\pi (4 - \frac 73 - \frac 14) = 2\pi (\frac {48}{12} - \frac {28}{12} - \frac {3}{12}) = 2\pi (\frac{48-28-3}{12}) = \\ \Large 2\pi (\frac{17}{12}) = \frac{17}{6}\pi \)

OpenStudy (anonymous):

Woah, I was wayy off

OpenStudy (anonymous):

Oh, I see where I messed up. The third fraction I put 12, as id I were multipying the whole thing by 12.

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (aum):

You are welcome.